# A Diffusion-based Xray2MRI Model: Generating Pseudo-MRI Volumes From one Single X-ray

### v1

@@ 3 Experimental settings
This section provides an overview of the experimental data, the preprocessing steps undertaken, and the specifics of the experimental procedures.
@@ 3.2.1 X-ray
As shown in Fig. 3, the knee joint (Fig. 3(b)) was identified from the plain radiograph (Fig. 3(a)) through the YOLOv2 learning model [31], with images sized of 299 $\times$ 299 $\times$ 1, which were resized to the size of 256$\times$256$\times$1, and image intensity distribution was normalized to [-1, 1].
(a) A standard knee plain radiograph
(b) An identified knee joint
Figure 3: A standard knee plain radiograph from the database and an identified knee joint highlighted in a red box 3(a). An identified knee joint 3(b).
@@ 3.2.2 MRI
Figure 4: Slice count distribution on each type of MRI sequences 4. Grey value distribution on each proportional position of slices within the T1-weighted MRI sequences 4.
@@ 3.3 Experimental details
The diffusion module utilized pre-trained weights from Zero-123 [32]. The pre-trained KOA classifier $\mathcal{C}$ and encoder $\mathcal{E}_{2}$ are sourced from the [33] and [34], respectively. During training, a base learning rate of 1e-06 was applied, with a lambda linear scheduler incorporating a warm-up period every 100 steps. The batch size was set to 64, with $T$ configured to 1,000 steps, and an Exponential Moving Average (EMA) [35] of 0.99 was used for the diffusion sampling process. The scale factor of the latent space was set as 0.2. The AdamW optimizer [36] was utilized for training over 5,000 epochs. We implemented our approach using PyTorch v1.8.1 [37] on Nvidia A100 graphics cards with 80 GB of memory.
@@ 4.1 Comparisons with State-of-the-Art Methods
In this section, the comparison is divided into two main parts: Visualization analysis and Quantitative indicator analysis.
4.1.1 Visualization analysis
Table 2 provides a comprehensive visual comparison of corresponding pseudo-MRI slices generated by various methods, using X-ray images as input. To facilitate a clear and focused analysis, we selected three representative stages of KOA: KL-0, indicating healthy knees; KL-2, reflecting early-stage KOA; and KL-4, corresponding to late-stage KOA. The compared methods include classic LDM [38], Cola-diff [20], X-diff [39], and our proposed approach. Each row displays the X-ray input on the far left, followed by the real MRI slice and the pseudo-MRI slices generated by the respective methods. To provide a more intuitive illustration of each method’s performance, difference maps generated using a Sobel operator filter are shown beneath each slice, highlighting pixel-wise deviations from the ground truth MRI by detecting edge differences. The visual results consistently demonstrate that our model produces pseudo-MRI slices that more closely resemble the ground truth, with significantly fewer and less pronounced deviations compared to other methods. Specifically, LDM suffers from the lack of constraints provided by the input X-ray. Without this conditional guidance, the model struggles to capture the detailed structural information of the knee joint, leading to great deviations from the ground truth MRI. On the other hand, compared with LDM, although Cola-Diff and X-Diff generate bone structures that are closer to the real MRI in terms of overall shape, they both lack a significant amount of fine details. This limitation prevents them from accurately capturing and reproducing the KOA symptoms. Moreover, it is noteworthy that it is particularly evident as the KL grade of the input X-ray increases. The progressive narrowing of the knee joint space and the formation of osteophytes are learned and generated more accurately by our model, which demonstrates that our method not only ensures the generation of high-quality MRI images but also maintains clinical relevance. However, the model’s predictions in areas outside the skeletal regions are less accurate due to the absence of critical information in X-ray images such as muscle and fat distribution, which will be discussed in in Section 4.3.3.
Table 2: Visualization of the generated pseudo-MRI slices
@@ 4.1.1 Visualization analysis
    To make it more intuitive and convenient, we only display here a single slice from each MRI sequence that best represents the complete knee joint.
4.1.2 Quantitative indicator analysis
Peak Signal-to-Noise Ratio (PSNR) is a measure used to assess the quality of reconstructed or compressed images compared to their original versions. The formula for PSNR is:
$\text{PSNR}=10\cdot\log_{10}\left(\frac{\text{MAX}_{I}^{2}}{\text{MSE}}\right)$ (18)
where $\text{MAX}_{I}$ is the maximum possible pixel value of the image. MSE stands for the Mean Squared Error, which calculates the average of the squared differences between the real and synthesized images:
$\text{MSE}=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}(I(i,j)-K(i,j))^{2}$ (19)
where $I(i,j)$ and $K(i,j)$ are the pixel values at position $(i,j)$ in the real and synthesized images, respectively. $m$ and $n$ are the width and height of the image.
On the other hand, the Structural Similarity Index (SSIM) is a widely used metric for evaluating the visual similarity between two images considering human visual perception by measuring changes in luminance, contrast, and structural information. However, it does not account for the specific regions of interest (ROI) that are critical in medical imaging. In knee MRI volumes, the most relevant diagnostic information often lies within cartilage, joints, and other areas affected by disease. Therefore, in addition to SSIM, we also adopt a specially designed indicator, the so-called Region-specific Structural Similarity Index (RSSIM), to represent the clinical evaluation of bone morphology. To do so, we utilize the Canny operator with a sigma of 3 to filter and extract bone edges, followed by the application of SSIM. Specifically, we focus on the lower third of the femoral edge and the upper third of the tibial edge to capture ROIs most likely to contain areas of potential pathology, such as joint space narrowing (JSN) and osteophytes. The typical SSIM is computed as:
$\text{SSIM}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (20)
where
$\begin{split}&\mu_{x}=\frac{1}{N}\sum_{i=1}^{N}x_{i},\quad\mu_{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}\\ &\sigma_{x}^{2}=\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\mu_{x})^{2},\quad\sigma_{y}^{2}=\frac{1}{N-1}\sum_{i=1}^{N}(y_{i}-\mu_{y})^{2}\\ &\sigma_{xy}=\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\mu_{x})(y_{i}-\mu_{y})\end{split}$ (21)
$\mu_{x}$, $\mu_{y}$, and $\sigma_{x}^{2}$, $\sigma_{y}^{2}$ are mean intensities and variances of the images $x$ and $y$, respectively. $\sigma_{xy}$ represents the covariance of $x$ and $y$. $C_{1}$ and $C_{2}$ serve as the stabilization constants to avoid division by zero.
@@ 4.1.2 Quantitative indicator analysis
4.2 Impact of the inference times
For the inference times $t$, in addition to adhering to the original depth setting ($d=t=50$), we also experimented with varying inference times. The goal was to explore the impact of increasing inference times on the quality and accuracy of the generated slices. For more clarity and conciseness, we here only focus on discussing the case where $t=500$. Table 3 illustrates the synthesized intermediate frames between the original adjacent images (i.e., starting and ending images). Three distinct representative stages were evaluated: first, the transition from predominantly muscle tissue to the initial appearance of bone; second, the shift from cartilage to the developing coronal knee joint; and finally, the full visualization of the complete coronal knee joint. In each of these stages, the synthesized intermediate frames display relatively smooth and continuous transitions between bone and soft tissue structures. Importantly, the model successfully maintains key anatomical details throughout the transition, including the integrity of the bones, cartilage, and joint spaces, which underscores the robustness and effectiveness of these interpolation frames.
Moreover, the Fourier frequency domain analysis and histogram correlation measurements were conducted. As shown in Fig. 6, the intermediate frames are almost consistent with the preceding and following frames in the low-frequency region, while they contain less high-frequency information, making the overall image appear smoother, which suggests that the intermediate frames are generated through interpolation. The average histogram correlations between the original adjacent images and their four intermediate frames are 0.5404, 0.9713, 0.9334, 0.8723, and 0.5248, in sequential order. The high correlations between the consecutive intermediate frames suggest that the intensity distribution remains nearly constant throughout these frames, indicating very little variation in the pixel intensities, which points to the fact that the intermediate frames are smooth transitional frames. In contrast, the relatively lower correlations between the real start and end frames and the closest intermediate frames highlight the greater difference in intensity distributions, which suggests that the interpolation algorithm focuses on generating a gradual and seamless transition between the start and end frames, prioritizing continuity over the introduction of new or abrupt changes.
(a) Starting image
(b) Frame #$\#1$
(c) Frame #$\#2$
(d) Frame #$\#3$
(e) Frame #$\#4$
(f) Ending image
Figure 6: The Fourier frequency domain distribution for the original adjacent images and the synthesized intermediate frames between them.
The above results demonstrate that increasing the inference times $t$ significantly enhances interpolation, showcasing its potential for improving temporal resolution. By generating additional intermediate frames from a smaller set of real frames, interpolation effectively offers more continuous data without lengthening the actual scan duration.
@@ 4.3 Discussion
In this paper, we introduce a novel diffusion-based Xray2MRI model that generates full pseudo-MRI volumes from a single X-ray image, representing one promising initial step toward cost-effective medical imaging solutions. The model leverages conditional inputs from X-rays, combined with target depth, KOA probability distribution, and image intensity distribution modules, to guide the synthesis process and ensure anatomically accurate MRI slice generation. Experimental results show that the generated pseudo-MRI sequences approximate real MRI scans, delivering high image quality while preserving critical clinical relevance. Additionally, by increasing the inference times, the model exhibits effective interpolation, resulting in smoother transitions between slices. Several points are to be discussed.
@@ 4.3.1 How it works?
In this approach, we move away from focusing on the distinct imaging principles and techniques of X-ray and MRI. Instead, we only concentrate on the mapping relationship between these two imaging modalities. Much like an extremely experienced senior radiologist who, when viewing an X-ray image, could to some extent mentally infer what the corresponding MRI scans might reveal, which is our proposed AI model’s aim. To this end, for the model’s training, vast amounts of high-quality paired X-ray and MRI, along with various data augmentation techniques, have significantly propelled this.
@@ 4.3.3 Strengths and limitations
This study has several notable strengths. Firstly, it introduces an innovative method for generating pseudo-MRI images from one single X-ray image, attempting to bridge the gap between these two imaging modalities using advanced AI technology. This approach represents a novel attempt in cases where MRI may not be readily available. Second, the study employs a 2D diffusion model, which not only reduces the computational resources required but also allows for more detailed control over the generated images, leading to a more interpretable and transparent reconstruction process. Finally, the study’s focus on clinically relevant metrics, including both traditional image quality assessments and pathology-specific evaluations, ensures that the generated pseudo-MRI images are not only visually similar to ground truth MRIs but also clinically relevant. This study has several limitations that warrant consideration. First, our model are highly dependent on the availability of large, high-quality paired datasets. The scarcity of such datasets, particularly in certain clinical scenarios or for rare diseases, may limit the generalizability of our approach. Second, although our model is able to approximate MRI features from X-ray inputs, it remains strongly uncertain whether these pseudo-MRI images can be reliably used for clinical diagnosis without the support of additional imaging modalities as the model cannot learn information that is not present in the input X-ray image, which is precisely the value that MRI imaging provides. Such limitation is clearly illustrated in Fig. 7. Finally, the computational resources required for training, despite being reduced by using a 2D model, high-performance hardware is still necessary, which presents a challenge for broader implementation in clinical settings.
@@
Figure 7: The X-ray 7(a) does not provide information about cartilage abnormalities, so the generated MRI 7(b) similarly lacks this detail, whereas such information is presented in the actual MRI 7(c).
@@ 4.3.4 Future work
More comprehensive information (e.g., BMI, age, gender, etc) could be effectively integrated with X-ray images to further enhance the generation performance. Moreover, to improve the model’s deployability, the use of knowledge distillation could be of interest.

### v2

@@ 3.2.1 X-ray
As shown in Fig. 3, the knee joint (Fig. 3) was identified from the plain radiograph (Fig. 3) through the YOLOv2 learning model [29], with images sized of 299 $\times$ 299 $\times$ 1, which were resized to the size of 256$\times$256$\times$1, and image intensity distribution was normalized to [-1, 1].
Figure 3: A standard knee plain radiograph and an identified knee joint highlighted in the red box 3. An identified knee joint 3.
@@ 3.2.2 MRI
Figure 4: Distribution of the number of slices for each type of MRI sequence 4. Grey value distribution at each proportional position of slices within the T1-weighted MRI sequences 4.
@@ 3.3 Experimental details
The diffusion module utilized pre-trained weights from Zero-123 [30]. The pre-trained KOA classifier $\mathcal{C}$ and encoder $\mathcal{E}_{2}$ were sourced from the [31] and [32], respectively. During the training, a base learning rate of 1e-06 was applied, with a lambda linear scheduler incorporating a warm-up period every 100 steps. The batch size was set to 64, with $T$ configured to 1,000 steps, and an Exponential Moving Average (EMA) [33] of 0.99 was used for the diffusion sampling process. The scale factor of the latent space was set as 0.2. The AdamW optimizer [34] was utilized for training over 5,000 epochs. We implemented our approach using PyTorch v1.8.1 [35] on Nvidia A100 graphics cards with 80 GB of memory.
@@ 4.1 Comparisons with State-of-the-Art Methods
In this section, the comparison is divided into two main parts: Qualitative visualization analysis and Quantitative metric-based analysis.
4.1.1 Qualitative visualization analysis
Table 1 provides a comprehensive visual comparison of corresponding pseudo-MRI slices generated by various methods, using X-ray images as input. To facilitate a clear and focused analysis, we selected three representative stages of KOA: G-0, indicating healthy knees; G-2, reflecting early-stage KOA; and G-4, corresponding to late-stage KOA. The compared methods include classic LDM [36], Cola-diff [18], X-diff [37], and our proposed approach. Each row displays the X-ray input on the far left, followed by the real MRI slice and the pseudo-MRI slices generated by the respective methods. To provide a more intuitive illustration of each method’s performance, difference maps generated using a Sobel operator filter are shown beneath each slice, highlighting pixel-wise deviations from the ground truth MRI by detecting edge differences. The visual results consistently show that our model produces pseudo-MRI slices that are closer to the ground truth, with significantly fewer and less pronounced deviations compared to other methods. Specifically, LDM suffers from the lack of constraints provided by the input X-ray. Without this conditional guidance, the model struggles to capture the detailed structural information of the knee joint, leading to great deviations from the ground truth MRI. On the other hand, compared with LDM, although Cola-Diff and X-Diff generate bone structures that are closer to the real MRI in terms of overall shape, they both lack a significant amount of fine details. This limitation prevents them from accurately capturing and reproducing the KOA symptoms. Moreover, it is noteworthy that it is particularly evident as the KOA grade of the input X-ray increases. The progressive narrowing of the knee joint space and the formation of osteophytes are learned and generated more accurately by our model, which demonstrates that our method not only ensures the generation of high-quality pseudo-MRI images but also maintains clinical relevance. However, the model’s predictions in areas outside the skeletal regions are less accurate due to the absence of critical information in X-ray images such as muscle and fat distribution, which will be discussed in Section 4.3.3.
Table 1: Visualization of a representative generated pseudo-MRI slice
@@ 4.1.1 Qualitative visualization analysis
    Here, to make it more intuitive and convenient, we only display a single slice from each MRI sequence that best represents the complete knee joint.
4.1.2 Quantitative metric-based analysis
@@
4.2 Impact of the number of inference steps
For the inference steps $s$, in addition to adhering to the original depth setting (i.e., $s=S=50$), we also experimented with varying inference steps (i.e., 30, 100, 200, 300, 400, and 500), to explore the impact of increasing the number of inference steps on the quality and accuracy of the generated slices. Fig. 6 shows the adjacent slice correlation coefficients under different inference steps. As can be seen, the red horizontal lines at 0.73 and 0.82 represent the mean adjacent slice correlation coefficients of the pseudo-MRI matching the volume of the original MRI (i.e., $s=S=50$) and that of the original MRI itself, respectively, with the latter serving as the upper benchmark for comparison. At 30 inference steps, the adjacent slice correlation starts at approximately 0.66, reflecting the impact of downsampling, which introduces inconsistencies and reduces the alignment between neighbouring slices. As the number of inference steps increases to 50, where the depth aligns with that of the original MRI volume, the correlation coefficient jumps sharply to around 0.73, suggesting a significant improvement in slice coherence. At 100 steps, the correlation continues to increase modestly. Beyond 100 steps, the improvement becomes more gradual as the number of inference steps increases, with diminishing returns, eventually stabilizing around 0.76 by 500 inference steps.
Figure 6: The adjacent slice correlation coefficients under different inference steps.
On the other hand, for the sake of clarity and conciseness, we limit our visualization to the interpolation at $s=500$. Table 2 illustrates the synthesized intermediate frames between the original adjacent images (i.e., starting and ending images). Two distinct representative stages were evaluated: The transition from predominantly muscle tissue to the initial appearance of bone, and the shift from cartilage to the developing coronal knee joint. In each of these stages, the synthesized intermediate frames display relatively smooth and continuous transitions between bone and soft tissue structures. Importantly, the model successfully maintains key anatomical details throughout the transition, including the integrity of the bones, cartilage, and joint spaces, which underscores the robustness and effectiveness of these interpolation frames.
@@ 4.2 Impact of the number of inference steps
The above results demonstrate that increasing the number of inference steps $s$ enhances interpolation, showcasing its potential for improving temporal resolution. By generating additional intermediate frames from a smaller set of real frames, interpolation effectively offers more continuous data without lengthening the actual scan duration. On the other hand, despite the improvements observed, the correlation coefficient from the interpolated pseudo-MRI fails to reach the 0.82 benchmark of the original MRI volume, even with the maximum 500 inference steps, which indicates that, while the interpolation process can enhance the consistency between adjacent slices, inherent limitations remain. These limitations arise from the inherent constraints of using X-rays as the sole source of information, which will be discussed in Section 4.3.3.
@@ 4.3 Discussion
In this paper, we introduce a novel diffusion-based Xray2MRI model that generates full pseudo-MRI volumes from a single X-ray image, representing one promising initial step toward cost-effective medical imaging solutions. The model leverages conditional inputs from X-rays, combined with target depth, KOA probability distribution, and image intensity distribution modules, to guide the synthesis process and ensure anatomically accurate MRI slice generation. Experimental results show that the generated pseudo-MRI sequences approximate real MRI scans, delivering high image quality while preserving critical clinical relevance. Additionally, by increasing the number of inference steps, the model exhibits effective interpolation, resulting in smoother transitions between slices. Several points are to be discussed.
@@ 4.3.1 How it works?
In this approach, we move away from focusing on the distinct imaging principles and techniques of X-ray and MRI. Instead, we only concentrate on the mapping relationship between these two imaging modalities. Much like an extremely experienced senior radiologist who, when viewing an X-ray image, could to some extent mentally infer what the corresponding MRI scans might reveal, which is our proposed AI model’s aim. To this end, for the model’s training, vast amounts of high-quality paired X-ray and MRI have significantly propelled this.
@@ 4.3.3 Strengths and limitations
This study has several notable strengths. Firstly, it introduces an innovative method for generating pseudo-MRI images from one single X-ray image, attempting to bridge the gap between these two imaging modalities using advanced AI technology. This approach represents a novel attempt in cases where MRI may not be readily available. Our experiments demonstrate that incorporating supplementary information beyond the X-ray images themselves, such as disease probability distributions, enhances the quality of the generated pseudo-MRI, which underscores the potential of integrating more diverse information sources in the future to produce more accurate pseudo-MRI volumes. Second, the study employs a 2D diffusion model, which not only reduces the computational resources required but also allows for more detailed control over the generated images, leading to a more interpretable and transparent reconstruction process. Finally, the study’s focus on clinically relevant metrics, including both traditional image quality assessments and pathology-specific evaluations, ensures that the generated pseudo-MRI images are not only visually similar to ground truth MRIs but also clinically relevant. This study has several limitations that warrant consideration. First, our model are highly dependent on the availability of large, high-quality paired datasets. The scarcity of such datasets, particularly in certain clinical scenarios or for rare diseases, may limit the generalizability of our approach. Second, although our model is able to approximate MRI features from X-ray inputs, it remains strongly uncertain whether these pseudo-MRI images can be reliably used for clinical diagnosis without the support of additional imaging modalities as the model cannot learn information that is not present in the input X-ray image, which is precisely the value that MRI imaging provides. Such limitation is clearly illustrated in Fig. 7. Finally, the computational resources required for training, despite being reduced by using a 2D model, high-performance hardware is still necessary, which presents a challenge for broader implementation in clinical settings.
@@
Figure 7: The X-ray 7(a) does not provide information about cartilage abnormalities, the generated pseudo-MRI 7(c) thus similarly lacks this detail, whereas such information is presented in the real MRI 7(b).
@@ 4.3.4 Future work
More comprehensive information (e.g., BMI, age, gender, disease type and location, etc) could be effectively integrated with X-ray images to further enhance the generation performance. Moreover, to improve the model’s deployability, the use of knowledge distillation could be of interest.